Opportunity cost, comparative advantage and gains from trade

Ali and Barb both like making bread and jam. In one day Ali can make 10 loafs of bread OR
20 jars of jam. In one day Barb can make 6 loafs of bread OR 36 jars of jam.


3a) What is Ali’s opportunity cost of making 1 loaf of bread? (Include a definition of opportunity cost).
3b) What is Barb’s opportunity cost of making one loaf of bread?
3c) If they both split their time evenly between bread making and jam making, how many loafs of bread
and jars of jam would they each have?
3d) If Ali only made bread and Barb only made jam, how many loafs of bread and jars of jam would they
have between them?
3e) If Ali only makes jam and Barb only makes bread, how many loafs of bread and jars of jam would they
have between them?
3f) Assume Ali and Barb specialize in their comparative advantage good, and then agree to trade 3 loafs
of bread for 12 jars of jam. How many loafs of bread and jars of jam do they each have now?

3g) Has specializing in their comparative advantage good made them better off overall? Explain your
answer fully, including explaining what is meant by ‘comparative advantage’.
Answer #1

Question 3

(a)

Opportunity cost of producing a good refers to the amount of other good that has to be sacrificed in order to produce one unit of given good.

Ali can produce either 10 loafs of bread or 20 jars of jam in one day.

Calculate the opportunity cost of making 1 loaf of bread -

Opportunity cost = Quantity of Jam produced in one day/Quantity of bread produced in one day

Opportunity cost = 20/10 = 2

Thus,

Ali's opportunity cost of making 1 loaf of bread is 2 jars of Jam.

(b)

Barb can produce either 6 loafs of bread or 36 jars of jam in one day.

Calculate the opportunity cost of making 1 loaf of bread -

Opportunity cost = Quantity of Jam produced in one day/Quantity of bread produced in one day

Opportunity cost = 36/6 = 6

Thus,

Barb's opportunity cost of making 1 loaf of bread is 6 jars of Jam.

(c)

In one day, Ali can make either 10 loafs of bread or 20 jars of jam.

So, if Ali split his time evenly between bread making and jam making then he will use half the day to make bread and half the day to make jam.

He would be able to produce 5 loafs of bread and 10 jars of jam.

In one day, Barb can make either 6 loafs of bread or 36 jars of jam.

So, if Barb split his time evenly between bread making and jam making then he will use half the day to make bread and half the day to make jam.

He would be able to produce 3 loafs of bread and 18 jars of jam.

(d)

In one day, Ali can make a maximum of 10 loafs of bread.

In one day, Barb can make a maximum of 36 jars of Jam.

Thus,

If Ali only made bread and Barb only made jam, they would have 10 loafs of bread and 36 jars of jam between them.